Structuring the Sell-or-Process-Further Decision

Jack's Lumber Yard receives 8,000 large trees each period that it subsequently processes into rough logs by stripping off the tree bark and leaves. Jack’s then must decide whether to sell its rough logs (for use in log cabin construction) at split-off or to process them further into refined lumber (for use in regular construction framing). Jack's normally sells logs for a per-unit price of $525. Alternately, each log can be processed further into 800 board feet of lumber at an additional cost of $0.05 per board foot. Also, lumber can be sold for $0.75 per board foot. (Note: One tree is equal to one rough log.)

Required:

1. What is the total contribution to income from selling the logs for log cabin construction?
$

2. What is the total contribution to income from processing the logs into lumber?
$
Answer #1

1. What is the total contribution to income from selling the logs for log cabin construction?

Contribution= 8000*$525= $4200000

2. What is the total contribution to income from processing the logs into lumber?

Further processing cost= 800*$0.05*8000= $320000

Sales price= 800*$0.75*8000= $4800000

Contribution= Sales price-Further processing cost

= $4800000-320000= $4480000